Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 16:24 Friday 17 March 2023 (24444) Print this report
Sensor correction and blending test.

I measured the performance of the new sensor correction and blending filters. Analysis will be reported later.

I wrote the blending filters lp_part1 and lp_part2 for LVDTL, and hp_part1 and hp_part2 for ACCL. Each blending filter had to be split in two multiplicative factors because their orders were very high.

The aim was to measure the ASDs of three different quantities when the suspension was swinging free. Namely, we wanted the readout of

  • The corrected LVDT, blended with the geophones.
  • The corrected LVDT only.
  • The  uncorrected LVDT.

In order to measure these quantities simultaneously, I did the following:

  1. Set the suspensionsion in SAFE state.
  2. Turned off the old blending filters (LP80_new and HP80_new) and turned on the new ones.
  3. Set the state to TWR_FLOAT, and then I paused Guardian.
  4. Turned on the sensor correction for LVDTL.
  5. Modified the matrices LVDT2EUL and LVDTALIGN in order to get the readout in L in the channel for T (see klog 22376).
  6. Ran  the measurement with diaggui.
  7. After the measurement, I reverted the changes to the matrices and made sure  the suspension went to LOCK_ACQUISITION state successfully.
  8. I letf the new filters on and modified the SDF in SAFE state.

In  the diaggui file, the relevant channels are as follows:

  • Corrected LVDT without blending: K1:VIS-BS_IP_BLEND_LVDTL_IN1.
  • Corrected LVDT blended with the geophone: K1:VIS-BS_IP_IDAMP_L_IN1.
  • Uncorrected LVDT: K1:VIS-BS_IP_BLEND_LVDTT_IN1 (as explained above, this is a T channel, arranged to provide information about the L channel).
  • Seismic motion along L, pre-filtered and in units of micrometers: K1:VIS-BS_IP_SENSCORR_L_IN1.

The location andname of the diaggui file is

  • /kagra/Dropbox/Subsystems/VIS/vis_commissioning/bs/data/
  • bs_ip_twr_float_04.xml

The new filters are written the file attached.

Non-image files attached to this report
Comments to this report:
fabian.arellano - 16:29 Friday 24 March 2023 (24533) Print this report

I did additional measurements to assess the performance of the sensor correction filter in IP-L. Analysis is pending.

Analysis of the previous measurements, that has not been reported in the klog, shows that the sensor correction filter is not performing as expected. At 0.2 Hz, attenuation of the seismic noise in the corrected sensor by a factor of 2 was achieved, whereas attenuation by a factor of 5 was expected. Therefore, I tried to find the origin of the discrepancy.

I did the necessary measurements to calculate again the intercalibration factors between the seismometer and IP-L and IP-T (per klogs 20873 and 22229). Although a proper calculation is pending, a cursory inspection of the data suggests the previously calculated values are just fine (see klog 22442). The data is in the directory /kagra/Dropbox/Subsystems/VIS/vis_commissioning/bs/inter_calibration/ in the file bs_ip_lt_intercalibration_with_seis.xml.

Then, I repeated the measurements referred to in klog 24444, but without any diagonalization in the LVDTALIGN matrix. The channel T was used for the uncorrected LVDT-L. The matrix was

[[ 1   , 0 ,  0  ],
 [ 1  ,  0 ,  0   ],
 [ 0  ,  0  ,  1  ]]

The data is in the file /kagra/Dropbox/Subsystems/VIS/vis_commissioning/bs/data/diaggui bs_ip_twr_float_05.xml.

For the last measurement, I used the filter banks for the blending filters in order to subtract the motion of the IP table measured by the geophones, from the corrected LVDT readout. In principle, this is the corrected LVDT noise. This quantity was at the channel  K1:VIS-BS_IP_IDAMP_L_IN1. The data is in file /kagra/Dropbox/Subsystems/VIS/vis_commissioning/bs/data/diaggui bs_ip_twr_float_06.xml.

Once I finished, I set all the LVDTALIGN matrix to whatever it was before (see klog 22376).

Search Help